An investor invested a total of $2,500 in two mutual funds. One fund earned a 9% profit while the other earned a 3% profit. If the investor's total profit was $159, how much was invested in each mutual fund?
the amount invested in the mutual fund that earned 9% is
the amount invested in the mutual fund that earned 3% is

Answer by ReadingBoosters(3246)

x+y = 2500
.09x + .03y = 159
...
x = 2500 - y
...
.09(2500-y) + .03y = 159
225 - .09y + .03y = 159
-.06y = 159-225
-.06y = -66
y = $1,100 at 3%
...
x = 2500 - 1100 = $1,400 at 9%
.....................
Have more questions?
Email: HomeworkHelpers@readingboosters.com
Delighted to help.
-Reading Boosters
Website: www.MyHomeworkAnswers.com
Wanting for others what we want for ourselves.